1/1

Adriaen (François) Valéry,

We gather together

tradukita de Theodore Baker

We gather together to ask the Lord’s blessing; He chastens and hastens His will to make known. The wicked oppressing now cease from distressing. Sing praises to His Name; He forgets not His own.

Beside us to guide us, our God with us joining, Ordaining, maintaining His kingdom divine; So from the beginning the fight we were winning; Thou, Lord, were at our side, all glory be Thine!

We all do extol Thee, Thou Leader triumphant, And pray that Thou still our Defender will be. Let Thy congregation escape tribulation; Thy Name be ever praised! O Lord, make us free!

Traduko de la Nederlanda poemo “Wilt heden nu treden” de Adriaen (François) Valéry (Adrianus Valerius, ∗1575 – †1620-01-27) en la Anglan de Theodore Baker (∗1851 – †1934).

Arg-528-1047 (2010-03-22 21:09:32)

Pri la tradukinto Theodore Baker vidu la retejon http: // www. hymntime. com/ tch/ bio/ b/ a/ k/ baker_ t. htm . Li angligis la tekston de la kanto laŭ la Germanigo farita de la aŭstro Eduard Kremser (1838 - 1914). Pri Eduard Kremser vidu: http: // www. hymntime. com/ tch/ bio/ k/ r/ kremser_e. htm .

http://www.poezio.net/version?poem-id=528&version-id=1047